a boat takes 4 hours to go 640km downstream. it takes the same to go 240 km upstream. find the speed of boat in still water and the speed of the stream.
Answer by richwmiller(17219)

r*t=d
with the current
(r+c)*4=640
(r+c)=640/4
r+c=160
against the current
(r-c)*4=240
(r-c)=240/4
r-c=60
r+c=160
add
r+r=60+160
2r=220
r=110
r=110 boat
110-c=60
c=110-60
c=50 current
check
(110+50)*4=640
160*4=640
640=640
ok
(r-c)*4=240
(110-50)*4=240
60*4=240
240=240
ok